Grade 12th passMechanics

A marble block of mass 2 kg lying on ice when given a velocity of 6 m/s is stopped by friction in 10s. Then the coefficient of friction is (consider g =10m/s)
(a) 0.02
(b) 0.03
(c) 0.06
(d) 0.01

Harshit Singh
5 Years agoGrade 12th pass

1 Answer

Pawan Prajapati
5 Years ago
Answer ::

u = 6 m/s
v = 0
t =10s
a = -f/m = – μmg/m = – μg = -10μ
Substituting values in v = u+at
0 = 6 – 10μ x 10
Therefore, μ = 0.06
Answer: (c) 0.06


Thanks